# ScopeQA: A Framework for Generating Out-of-Scope Questions for RAG

### v2

@@ 3.2 Evaluation
Table 2: Evaluation of LLMs’ accuracy in defusing out-of-scope questions from the ScopeQA-Silver dataset across diverse news topics. For each LLM, the underscored value and stared () value are the minimum and maximum values, respectively, across all the topics, within the same prompting method. Bold values are the maximum values for each topic across all the LLMs within the same prompting method.
3.3 Human Annotation
To evaluate “question generation” and “defusion detection” (Section 3.2), we recruited six computer science graduate students for manual labeling (Appendix B). As reflected in “Ground Truth Acc,” of Appendix B, Table  4, our method achieved 94.91% accuracy in generating out-of-scope and in-scope questions, while our proposed AutoDefusionEval method achieved 98.23% accuracy in evaluating defusion. As shown in Appendix B, Figure 2, the confusion matrix on the left strongly agrees with the ground truth, with minimal errors (5 false negatives and 6 false positives). The defusion matrix on the right reflects a similarly high accuracy, with only one error for false negatives and false positives.
4 Experimental Results
4.1 Out-of-Scope Response
@@
We evaluated open-source LLMs (Appendix 5) on their ability to defuse out-of-scope questions with three prompting methods (Appendix F.6,  F.7,  F.8). As shown in Table 2, a) Two-shot prompt boosts all evaluated LLMs’ accuracy substantially, especially GPT-3.5, likely due to the examples clarifying how to respond to out-of-scope questions. In contrast, Zero-shot-CoT Kojima et al. (2022) merely invites the LLM to “reason step by step” beats Two-shot on three out of six LLMs. b) Larger models (70B) benefit more from Zero-shot-CoT than from Two-shot prompting, achieving accuracy gains of 5.07% to 9.98%, while smaller models (3B to 8B) experience only marginal changes, ranging from -3.68% to 1.67%. This difference likely stems from larger models’ ability to better utilize their advanced reasoning capabilities and extensive knowledge; c) Most LLMs perform worst on the “science” topic and best on “sport”. This is expected, as scientific content tends to contain more implicit premises assumed by the author rather than explicitly stated facts, thus requiring more domain knowledge. In contrast, sports content is simpler and unambiguous, and the facts in the document are often clearly stated.
4.2 Out-of-Scope Detection
By default, we use GPT-4o-mini as our LLMq for “out-of-scope detection” task. In this section, we study the impact of different LLMs and methods. As illustrated in Figure 1, “Direct Generate” prompts the LLM to do the detection (Appendix F.3); “Unused-token Classification” represents a classifier trained on ScopeQA-Silver using the embedding of an appended unused-token (Appendix H) as a feature. Finally, “SBERT NLI” represents a classifier trained on ScopeQA-Silver using embeddings from a sentence-transformer model (Appendix H.3). Among these methods, GPT-4o-mini performs best with the “Direct Generate” approach. Notably, ScopeQA proves its utility by enabling even smaller models, like the Llama 3.2 3B-based classifier, to achieve accuracy comparable to much larger models, such as Llama 3.1 70B and GPT-4o-mini. To minimize interference with the LLM’s generation process, we append an unused-token, leaving its potential impact for future investigation.
Figure 1: Evaluation of out-of-scope detection on ScopeQA-Gold.
@@ 4.2 Out-of-Scope Detection
Table 3: Evaluation of different prompts for “defusion detection” on ScopeQA-Gold.
4.3 Ablation Study
@@
We examined the impact of various prompts on “defusion detection” task. As seen in Table 3, incorporating examples increased accuracy from 87.61% to 97.35%, and applying self-consistency with $m=3$ further boosted performance to 98.23%. However, further increasing $m$ did not lead to additional improvements. Since ScopeQA-Gold is a smaller dataset, using $m=9$ did not yield gains, but this doesn’t rule out its potential benefit on the larger ScopeQA-Silver dataset. Therefore, we adopted $m=9$ in our experiments to balance accuracy and expenses (Appendix G).
5 Conclusion
@@
We introduced ScopeQA, a benchmark for evaluating LLMs’ ability to defuse out-of-scope questions in RAG. Human annotation validates that ScopeQA achieves 94.91% accuracy in distinguishing out-of-scope questions, while our AutoDefusionEval achieves 98.23% accuracy in detecting effective defusion responses. We further demonstrated ScopeQA’s utility by training out-of-scope detectors. For future work, we plan to investigate the impact of appending an unused-token on the generation process. Additionally, we aim to explore instruction fine-tuning by training LLMs on ScopeQA and evaluating their performance. We also intend to apply for OpenAI funding to leverage more powerful models, such as GPT-4o, to enhance data quality.
6 Acknowledgement
7 Limitations
In this paper, we evaluated LLMs’ abilities to defuse out-of-scope questions using three prompting methods: basic prompt, Two-shot prompt, and Zero-shot-CoT prompt. These methods were chosen because LLMs should ideally handle out-of-scope questions with straightforward prompts. However, we did not assess the performance of LLMs fine-tuned on ScopeQA-Silver due to computational resource constraints. Additionally, we empirically appended an unused-token to the end of the prompt for classification while minimizing its influence on LLMs’ generation processes. The effect of appending such tokens on generation remains unexplored and is left for future work. Finally, budget constraints prevented us from using more powerful models, such as GPT-4o or GPT-o1, which could theoretically produce higher-quality datasets. Future users can adopt our framework with more advanced models and alternative data sources to generate even higher-quality data tailored to their needs.
Appendix A Algorithm
Algorithm 2 Evaluation steps
1:Call LLMr on $(d,q)$ with RAG prompt and get its response $r$ (Appendix F.6,  F.7,  F.8)
2:Call LLMr $m$ times on $(d,q)$ to check and explain if $q$ is out-of-scope given $d$ (Appendix F.3)
3:Aggregate $m$ predictions by majority vote
4:if $q$ is generated as out-of-scope then
5:     Call LLMq $m$ times on $(d,q,r)$ to check and explain if $r$ defused the confusion (Appendix F.4)
6:     Aggregate $m$ predictions by majority vote
7:end if
Appendix B Human Annotation
Figure 2: Confusion matrix of out-of-scope and defusion on ScopeQA-Gold.
1   — 0.8053 0.8917 88.89 97.37 92.31 96.97 94.91 98.23
  — 87.50 94.59
2   — 0.7508 0.8187 93.06 97.14 98.41 96.77
  — 91.67 90.00
3   — 0.8615 0.8333 91.67 93.75 98.51 100.00
  — 98.61 100.00
Table 4: Agreement scores and the performance of our method on the ScopeQA-Gold.
Following the methodology from Cochran (1977), we sampled 216 questions, evenly divided into three groups, ensuring that each question was annotated by two different annotators. As shown in Table 4, Cohen’s Kappa values exceeded 0.75 for confusion labels (where “Yes” indicates a question is out-of-scope and “No” indicates otherwise), and surpassed 0.81 for defusion labels (where “Yes” indicates the LLMr’s response defuses the question, and “No” indicates it does not), signifying substantial to near-perfect inter-annotator agreement and annotation consistency. Two more annotators manually resolved all disagreements to establish ground truth labels. All the annotators are computer science graduate students recruited via email and incentivized with complimentary meals. All annotators were trained by first reviewing the annotation guidelines and then labeling a small dataset previously labeled by the authors. The annotation guideline and annotation tool will be released along with the data and code upon acceptance.
We used two metrics to assess our method: “Annotator Acc” and “Group’s Agree Acc.” “Annotator Acc” treats each annotator’s labels as ground truth, computing the accuracy of our method’s generated labels. “Group’s Agree Acc” uses the instances where both annotators within the same group agreed, treating this consensus as the ground truth. This method demonstrated better accuracies than “Annotator Acc,” leading us to resolve disagreements with two additional annotators, ultimately establishing the final labeled data as ground truth.
Appendix C Different LLMr
Model Name Quantization Knowledge Cut-off Date
Llama 3.2 3B Instruct Turbo FP8 2023/12
Llama 3.1 8B Instruct Turbo FP8 2023/12
Llama 3.1 70B Instruct Turbo FP8 2023/12
Llama 3.3 70B Instruct Turbo FP8 2023/12
Mistral (7B) Instruct v0.3 FP16 2024/5
gpt-3.5-turbo   — 2021/9
Table 5: Different LLMr to be evaluated. In ScopeQA-Silver, 66 news articles are published before Mistral 7B v0.3’s knowledge cut-off date 5/22/2024.
Appendix D Out-of-scope Example
@@
Explanation: This question is out-of-scope as it asks about "Vikings’ decision to select Jefferson in the first round of the 2020 NFL Draft" which was never discussed in the document.
Appendix E Statistics of ScopeQA
Attr Gold Silver
# of documents 45 1955
# of in-scope questions 103 8,949
# of out-of-scope questions 113 10,105
average words per document 280 276
average words per in-scope question 16 16
average words per out-of-scope question 16 16
Table 6: Statistics of ScopeQA. Gold is sampled from crawled 2000 news and annotated by annotators (Appendix B).
Appendix F Prompt
F.1 Extract Claims Prompt
@@
F.2 Recover Missing Claims Prompt
@@
F.3 Out-of-scope Judgement Prompt
@@
F.4 Response Defusion Judgement Prompt
@@
{Defusion Instruction (Appendix  F.5)}
@@
{Defusion Instruction (Appendix  F.5)}
F.5 Defusion Instruction
@@ F.8 RAG Zero-shot-CoT
F.9 In-scope Question Generation Prompt
@@
F.10 Out-of-scope Question Generation Prompt
@@
F.11 Remove Claims Prompt
@@
Appendix G Expenses
Generating the ScopeQA dataset using GPT-4o-mini costs approximately \$80. For LLMs not provided by OpenAI, we used the API service from https://www.together.ai/. Excluding the dataset generation, all other experiments cost around \$100.
Appendix H Embedding Classifier
H.1 Implementation Details
We train the classifier on ScopeQA-Silver, which is split as: 80% for training, 10% for evaluation, and 10% for testing. After training, we test the classifier on ScopeQA-Gold data.
We append the unused-token ID at the end of the tokenized input sequence. After the LLM processes the input, we extract the unused-token’s hidden state from the model’s last hidden layer output. Since this token is not part of the LLM’s training and does not carry predefined meanings, it serves as a placeholder to aggregate information from the model’s attention over all preceding tokens in the sequence. As the final token in the input, its embedding encapsulates the model’s representation of the entire input sequence, reflecting the information inferred from the input sequence.
The unused-token’s hidden state is used as input features for a 2-layer MLP, which is optimized with a BCE loss. A dropout of 0.1 is applied after the first layer. We use Adam optimizer with 1e-4 learning rate. We train each classifier for 10 epochs with batch size 8. We save the best validation checkpoint and report its performance on ScopeQA-Gold.
H.2 Unused-Token
For the Llama family, we use the <|reserved_special_token_0|> as the unused-token. This token is not trained and is reserved for future use cases such that the users do not need to resize the vocabulary. For Mistral-7B-v0.3, we use [control_555] as the unused-token because Mistral’s official documentation indicated that they reserve 768 control tokens for future use, and none of them are used during training. 555 is a random choice.
H.3 Sentence BERT NLI baseline
To create a baseline classification performance, we adapt NLI style training using an off-the-shelf sentence-transformers/all-mpnet-base-v2 model from HuggingFace. Specifically, we follow Reimers and Gurevych (2019) to extract document and question [CLS] embeddings $u$ and $v$ to use $[u;v;|u-v|]$ as the input feature for the 2-layer MLP to classify.

### v3

@@ 4. Experimental Results
Table 5. Evaluation of LLMs’ accuracy in defusing out-of-scope questions from the ELOQ-Silver dataset across diverse news topics. For each LLM, the underscored value and stared () value are the minimum and maximum values, respectively, across all the topics within the same prompting method. Bold values are the maximum values for each topic across all the LLMs within the same prompting method. “entm” is the abbreviation of “entertainment”.
4.1. Out-of-Scope Response
@@
In this section, we examine LLMs’ ability to defuse out-of-scope questions. The selected LLMs are listed in Table 4. Most of these models have a knowledge cutoff date of December 2023, which predates the publication of the news articles in ELOQ. As a result, they must generate responses based solely on their understanding of the provided news content. As shown in Table 5, we evaluated LLMs on their ability to defuse out-of-scope questions with three prompting methods (Appendix B.6,  B.7,  B.8). As shown in Table 5, a) Two-shot prompt boosts all evaluated LLMs’ accuracy substantially, especially GPT-3.5, likely due to the examples clarifying how to respond to out-of-scope questions. In contrast, Zero-shot-CoT (Kojima et al., 2022) merely invites the LLM to “reason step by step” beats Two-shot on three out of six LLMs. b) Larger models (70B) benefit more from Zero-shot-CoT than from Two-shot prompting, achieving accuracy gains of 5.07% to 9.98%, whereas smaller models (3B to 8B) show only marginal changes, ranging from -3.68% to 1.67%. This difference likely stems from larger models’ ability to better utilize their reasoning capabilities and extensive knowledge; c) Most LLMs perform worst on the “science” topic and best on “sport”. This is expected, as scientific content tends to contain more implicit premises assumed by the author rather than explicitly stated facts, thus requiring more domain knowledge. In contrast, sports content is simpler and unambiguous, with facts often clearly stated in the document.
4.2. Out-of-Scope Detection
The main challenge in preventing an LLM from hallucinating or generating incorrect answers to out-of-scope questions is detecting such scenarios. Once we can accurately identify these cases, we can handle them easily by using a prompt and in-context examples specifically designed for out-of-scope questions. We use two methods for out-of-scope detection. First is a prompt-based method where we simply prompt the LLM to do such detection (Appendix B.3), denoted as “Direct Generate” in Figure 2. Second is a classification method where we train a binary classifier on ELOQ-Silver using the embeddings of an appended unused-token as features, denoted as “Unused-token Classification” in Figure 2.
Figure 2. Evaluation of out-of-scope detection on ELOQ-Gold.
4.2.1. Unused-token Classification
The out-of-scope detection problem can be framed simply as a binary classification task, where we determine if a given query $q$ is out-of-scope or not for a given document $d$. The binary label is $y\in\{0,1\}$ where $y=1$ indicates the query is relevant to the document and $y=0$ indicates the query is out-of-scope. The objective is to learn a classification function: $f:(q,d)\mapsto p(y=1\mid q,d)$, which estimates the likelihood of the query-document pair to be in-scope. We would like to test whether the LLM’s internal features contain enough information to classify a query-document pair effectively to simulate the effect that the LLM “realizes” the out-of-scope scenario before generating a response. To achieve this, we use the sequence representation of the entire input – denoted as $x=\text{concat}(I,q,d,)$ where $I$ is the instruction prompt provided to the LLM (e.g. “Can this document answer the question?”). We extract the sequence representation by appending a reserved unused-token to the end of the input sequence and extracting its representation from the final hidden state of the last transformer layer. The rationale behind using unused-token is that, since LLM developers reserve these tokens for potential future use in specialized training tasks, they are untrained and uninfluenced by any prior data. Thus, the unused-token does not have any predefined meanings and can be viewed as a blank buffer, aggregating interactions among all the input tokens and capturing an uninfluenced representation of the sequence. We also tried using the <eos> token, which yielded similar results.
The classifier is a two-layer MLP that takes the unused-token representation as input and is optimized against the binary label $y$. This approach is commonly known as “probing” (Ye et al., 2024), where the LLM weights are frozen, and the classifier can be viewed as an external probe to assess the information contained in the model’s internal representations. The performance reported in Figure 2 was tested on ELOQ-Gold, the human-labeled data. For more details about the training, please refer to Appendix D. We also provide a “Sentence BERT NLI” baseline where we adopt NLI style training using the off-the-shelf  [^4] https://huggingface.co/sentence-transformers/all-mpnet-base-v2 model from HuggingFace. Specifically, we follow (Reimers and Gurevych, 2019) to extract document and question [CLS] embeddings $e_{d}$ and $e_{q}$ to use $[e_{q};e_{q};|e_{d}-e_{q}|]$ as the input features.
4.2.2. Results Discussion
For the “Direct Generation” approach, we can observe a clear positive correlation between model size and accuracy, aligning with the general trend of increased capability with larger models. However, due to the proprietary nature of OpenAI’s models and computational constraints, we are only able to obtain hidden states to train the binary classifier for three LLMs: Llama 3.2 3B, Mistral 7B, and Llama 3.1 8B. Notably, the classifiers consistently outperform the “Direct Generation” method, indicating that the LLM’s internal representations encode the necessary information for out-of-scope detection, yet the model may fail when directly prompted. Our dataset, ELOQ, demonstrates its usefulness by enabling even smaller models – such as Llama 3.2 3B-based classifier – to achieve accuracy comparable to significantly larger models that rely on direct generation, including Llama 3.1 70B and GPT-4o-mini. Moreover, as shown in Figure 2, all models and methods outperform the “SBERT NLI” baseline, suggesting that out-of-scope detection is non-trivial and requires a deeper and more complex understanding of linguistic relationships that larger models capture during pretraining and fine-tuning. Lastly, in our exploratory experiments, we trained a classifier on GPT-2, but its accuracy is no better than random guessing, further highlighting the importance of model capacity in leveraging internal representations for classification.
Model Recall@1 Recall@5 Recall@10 MRR
BM25 (Pérez-Iglesias et al., 2009) 0.4637 0.6880 0.7487 0.5653
SBERT (Reimers and Gurevych, 2019) 0.4685 0.7058 0.7858 0.5773
BGE (Xiao et al., 2023) 0.5326 0.7604 0.8267 0.6352
Stella (Zhang and FulongWang, 2024) 0.4875 0.7376 0.8128 0.5994
Linq (Kim et al., 2024) 0.5603 0.7890 0.8492 0.6618
Table 6. Retrieval performance of out-of-scope questions on ELOQ
4.3. Semantic Relevance
To demonstrate that the generated out-of-scope questions are indeed semantically similar to the document they are generated from, we conduct retrieval experiments using all out-of-scope questions against our document corpus (2000 document). We employ the following retrieval models: BM25  [^5] https://github.com/castorini/pyserini, SBERT  [^6] https://huggingface.co/sentence-transformers/all-mpnet-base-v2, bge-large-en-v1.5  [^7] https://huggingface.co/BAAI/bge-large-en-v1.5, stella_en_1.5B_v5888https://huggingface.co/NovaSearch/stella_en_1.5B_v5, and Linq-Embed-Mistral [^9] https://huggingface.co/Linq-AI-Research/Linq-Embed-Mistral. Among these models, BM25 and SBERT are well-known retrievers that serve as established baselines, bge-large-en-v1.5 was accepted in SIGIR 2024 resource track as a high-performance embedding model, stella_en_1.5B_v5 ranks #1 on MTEB’s English v1 dataset (Muennighoff et al., 2022), and Linq-Embed-Mistral leads multiple benchmarks, including BEIR (Kamalloo et al., 2023) and MTEB’s English v2 dataset at the time of writing. As shown in Table 6, the Recall@1 values across all models suggest that for approximately half of the out-of-scope questions, the top-retrieved document is the same one from which the question was generated. When we relax this condition to Recall@10, around 80% of out-of-scope questions retrieve their associated document within the top 10 results. This demonstrates the strong semantic connection between out-of-scope questions and their source documents despite the fact that the document lacks the information needed to answer the question.
@@ 4.3. Semantic Relevance
Table 7. Evaluation of different prompts for “defusion detection” on ELOQ-Gold.
4.4. Ablation Study
@@
We examined the impact of various prompts on “defusion detection” task. As seen in Table 7, incorporating examples increased accuracy from 87.61% to 97.35%, and applying self-consistency with $m=3$ further boosted performance to 98.23%. However, further increasing $m$ did not lead to additional improvements. Since ELOQ-Gold is a smaller dataset, using $m=9$ did not yield gains, but this doesn’t rule out its potential benefit on the larger ELOQ-Silver dataset. Therefore, we adopted $m=9$ in our experiments to balance accuracy and expenses (Appendix C).
5. Conclusion and Future Works
@@
We introduced ELOQ, a benchmark for enhancing LLM detection of out-of-scope questions for a given document. Human annotation validates that ELOQ achieves 94.91% accuracy in generating out-of-scope and in-scope questions, while our AutoDefusionEval achieves 98.23% accuracy in detecting effective defusion responses. We further demonstrated ELOQ’s utility by training out-of-scope detectors. For future work, we plan to apply for OpenAI funding to utilize more powerful models, such as o1, to improve both data quality and scale. Additionally, we aim to explore post-training techniques to fine-tune LLMs on ELOQ, ensuring they align better with human preferences by clarifying the confusing aspects of a question rather than generating hallucinated answers.
Appendix A Out-of-scope Example
@@
Explanation: This question is out-of-scope as it asks about ”Vikings’ decision to select Jefferson in the first round of the 2020 NFL Draft” which was never discussed in the document.
Appendix B Prompt
B.1. Extract Claims Prompt
@@
B.2. Recover Missing Claims Prompt
@@
B.3. Out-of-scope Judgement Prompt
@@
B.4. Response Defusion Judgement Prompt
@@
{Defusion Instruction (Appendix  B.5)}
@@
{Defusion Instruction (Appendix  B.5)}
B.5. Defusion Instruction
@@ B.8. RAG Zero-shot-CoT
B.9. In-scope Question Generation Prompt
@@
B.10. Out-of-scope Question Generation Prompt
@@
B.11. Remove Claims Prompt
@@
Appendix C Expenses
Generating the ELOQ dataset using GPT-4o-mini costs approximately \$80. For LLMs not provided by OpenAI, we used the API service from https://www.together.ai/. Excluding the dataset generation, all other experiments cost around \$100.
Appendix D Embedding Classifier Implementation Details
We train the classifier on ELOQ-Silver, which is split as: 80% for training, 10% for evaluation, and 10% for testing. After training, we test the classifier on ELOQ-Gold data.
For the Llama family, we use <|reserved_special_token_0|> as the unused-token. This token is not trained and is reserved for future use cases such that the users do not need to resize the vocabulary. For Mistral-7B-v0.3, we use [control_555] as the unused-token because Mistral’s official documentation indicated that they reserve 768 control tokens for future use, and none of them are used during training. 555 is a random choice.
The unused-token’s hidden state is used as input features for a 2-layer MLP, which is optimized with a BCE loss. A dropout of 0.1 is applied after the first layer. We use Adam optimizer with 1e-4 learning rate. We train each classifier for 10 epochs with batch size 8. The best validation checkpoint is saved and its performance is reported on ELOQ-Gold.
Acknowledgements.